The question in the book asks: Use substitution to solve the linear system.
0.25x - 1.25y = 10.25
x - 5y = 20
My x and y's both cancel out or equal zero. I don't understand.
Thanks

Answer by babynarah2007(12)

The question in the book asks: Use substitution to solve the linear system.
0.25x - 1.25y = 10.25
x - 5y = 20
My x and y's both cancel out or equal zero. I don't understand.
Thanks
0.25x - 1.25y = 10.25
x = 10.25/0.25 + 1.25y/0.25
x = 41 + 5y
we can define this equation : x - 5y = 41 (not)x - 5y = 20
so, 0.25(41+5y)-1.25y = 10.25
10.25 + 1.25y - 1.25y = 10.25

Both X and Y are null set, because they were cancelled out.